# Efficiently Ranking Software Variants with Minimal Benchmarks

### v1

@@ 6.1. Experimental details
We implemented our methods in Python; an archive of the code is available as supplementary material along with all additional figures and tables. We will report Kendall’s $\tau$ as mentioned previously (see section 3). All experiments were run on a single core on Intel Xeon Gold 6130 (16 cores, 2017), and our algorithms returned the best result found so far. Working with multiple cores can speed up the results almost linearly, and better results may be achieved with a longer timeout on a single core, since all methods except MILP are anytime algorithms. Since our methods are stochastic, we conducted our experiments on 10 different seeds and reported mean performance along with the confidence interval 95% and performed statistical tests when necessary. The timeouts were 50 minutes per task for all datasets. As a reminder, all methods use divide-and-conquer and iterative solving except MILP.
@@ 6.2. Datasets Used
Code Generation: HumanEval (Chen et al., 2021; Liu et al., 2023b)
@@
First, we looked at the HumanEval benchmark (Chen et al., 2021) and also considered the HumanEval Plus (Liu et al., 2023b) version with additional tests to correctly test generated code; furthermore, in the repository, the authors provide the samples generated by numerous models. We took these models as variants and the HumanEval tasks as tests, which gives us a matrix of zeros and ones depending on the pass $k$ score used. We took pass1 for both the classic version and the plus versions, giving us 2 benchmarks. These datasets contain 120 variants and 164 instances. The cost for these benchmarks is 1 for all tests, as no cost was provided in the original data.
Algorithm Selection: ASlib (Bischl et al., 2016)
@@
Second, we looked at comparisons of algorithm performance. There are a large number of scenarios in the ASlib benchmark (Bischl et al., 2016), and we selected only scenarios where there were at least 8 variants. In fact, when too few variants are available, it is not as relevant. When multiple versions of the same data are available, we only keep the one with the most data. This produced 23 benchmarks with one metric to optimize, and the mean running time of the test was used as a cost; this means that these are instances of the WRTSM problem. The idea of using cost is to use the fact that some more complicated instances can provide more information but at the cost of more runtime than, for example, 10 smaller tests; therefore, not adding weights would be an unfair comparison of tests. This leads to a range of diverse benchmarks, from answer set programming, constraint solving, SAT, and others, with fewer than 100 variants and fewer than 500 tests. The number of variants is usually in the range of 8-20. The cost in these benchmarks is the mean running time of the tests in all variants.
Configurable Systems (Lesoil et al., 2024; Alves Pereira et al., 2021; Kaltenecker et al., 2020; Lesoil et al., 2023)
@@
Third, we looked at datasets that measure the performance of configurable systems (Lesoil et al., 2024; Alves Pereira et al., 2021; Kaltenecker et al., 2020; Lesoil et al., 2023). Specifically, the dataset contains 8 unique benchmarks, but multiple metrics are measured; therefore, we generated variants of the benchmarks with an increasing number of metrics, but we limited ourselves to 4 variants of the same benchmark. This adds a pool of 19 additional benchmarks with up to 200 variants and thousands of tests. For sqlite we made only 4 versions despite 15 different output metrics. The cost in these benchmarks is the mean running time of the tests across all variants when available; if time is a metric, then it is also used as cost. For example, gcc-ctime has ctime as a cost and metric, whereas lingeling-conflicts has cost 1 for all tests since no time is in the metrics.
Program Repair: RepairBench (Silva and Monperrus, 2024)
@@
Finally, we looked at the performance of LLM in automatic program repair with RepairBench (Silva and Monperrus, 2024), which contained two benchmarks at the time: gitbugjava (90 tests, 23 variants) and defects4j (484 tests, 22 variants). A variant here is an LLM. Then, three versions of each benchmark were created based on the release dates of the LLMs. The first cut-off date is Oct. 2024 (9 variants left), and the second one is Jan. 2025 (17 variants left). Adding a final six additional benchmarks. The cost for these benchmarks is 1 for all tests, since the cost was global and not on a test basis.
@@ 6.3. RQ1: How many tests can be eliminated while maintaining the same ranking?
In order to tackle this question, we will run our method on all datasets targeting no ranking error. The results are in LABEL:table:gain. For 17 of the benchmarks, no cost reduction was found, which means that for 33/50 we found a cost reduction. Finding no cost reduction is interesting because it means that all test instances are relevant for the ranking of the variants. From another point of view, there are a number of benchmarks where the cost reduction reaches near 1, so that means that almost all tests are useless to rank the variants; only a few are relevant. Looking more specifically at the data, BISS generally outperforms other approaches, notable contenders are RANDOM and MILP. This is why in future experiments we only consider MILP, BISS and RANDOM. It may surprise the reader that MILP is not the best performing solution since, by definition, it finds an optimal solution; this is because we cannot encode the full problem in MILP as such MILP solves a more constrained problem. As was explained previously, MILP indeed finds an optimal solution to the more constrained problem, which is not always the best solution to the true problem.

### v2

@@ 6.1. Experimental details
We implemented our methods in Python; an archive of the code is available as supplementary material along with all additional figures and tables. We will report Kendall’s $\tau$ as mentioned previously (see section 3). There is no ground-truth data for the minimal benchmark, and it is prohibitively expensive to calculate. Our empirical evaluation focuses on the absolute reduction in the benchmarks and the relative ranking of the different algorithms considered. All experiments were run on a single core on Intel Xeon Gold 6130 (16 cores, 2017), and our algorithms returned the best result found so far. Working with multiple cores can speed up the results almost linearly, and better results may be achieved with a longer timeout on a single core, since all methods except MILP are anytime algorithms. Since our methods are stochastic, we conducted our experiments on 10 different seeds and performed statistical tests when necessary. The timeouts were 50 minutes per task for all datasets. As a reminder, all methods use divide-and-conquer and iterative solving except MILP. In our experiments, there was no tie in the ranks of the variants.
One challenge acquired is that while theoretically, our divide and conquer should guarantee that the solution has the desired Kendall value, that is not the case. When two solutions $s_{1}$, $s_{2}$ for the tests suites $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are merged, theoretically if they give perfect ranking respectively, then using $s=s_{1}\cup s_{2}$ for $\mathcal{T}=\mathcal{T}_{1}\cup\mathcal{T}_{2}$ should give perfect ranking but because of the objective framed in the linear regression this does not happen and a worse result is obtained.
@@ 6.2. Datasets Used
Code Generation: HumanEval (chen2021codex; evalplus)
@@
First, we looked at the HumanEval benchmark (chen2021codex) and also considered the HumanEval Plus (evalplus) version with additional tests to correctly test generated code; furthermore, in the repository, the authors provide the samples generated by numerous models. We took these models as variants and the HumanEval tasks as tests, which gives us a matrix of zeros and ones depending on the pass $k$ score used. We took pass1 for both the classic version and the plus versions, giving us 2 benchmarks. These datasets contain 120 variants and 164 instances. The cost for these benchmarks is 1 for all tests, as no cost was provided in the original data.
Algorithm Selection: ASlib (bischl_aslib_2016)
@@
Second, we looked at comparisons of algorithm performance. There are a large number of scenarios in the ASlib benchmark (bischl_aslib_2016), and we selected only scenarios where there were at least 8 variants. In fact, when too few variants are available, it is not as relevant. When multiple versions of the same data are available, we only keep the one with the most data. This produced 23 benchmarks with one metric to optimize, and the mean running time of the test was used as a cost; this means that these are instances of the WRTSM problem. The idea of using cost is to use the fact that some more complicated instances can provide more information but at the cost of more runtime than, for example, 10 smaller tests; therefore, not adding weights would be an unfair comparison of tests. This leads to a range of diverse benchmarks, from answer set programming, constraint solving, SAT, and others, with fewer than 100 variants and fewer than 500 tests. The number of variants is usually in the range of 8-20.
Configurable Systems (LesoilSGATBJ24; alvespereira:hal-02148791; kaltenecker2020interplay; lesoil2023)
@@
Third, we looked at datasets that measure the performance of configurable systems (LesoilSGATBJ24; alvespereira:hal-02148791; kaltenecker2020interplay; lesoil2023). Specifically, the dataset contains 8 unique benchmarks, but multiple metrics are measured; therefore, we generated variants of the benchmarks with an increasing number of metrics, but we limited ourselves to 4 variants of the same benchmark. This adds a pool of 19 additional benchmarks with up to 200 variants and thousands of tests. For sqlite we made only 4 versions despite 15 different output metrics. The cost in these benchmarks is the mean running time of the tests across all variants when available; if time is a metric, then it is used as cost. For example, gcc-ctime has ctime as a cost and metric, whereas lingeling-conflicts has cost 1 for all tests since no time is in the metrics.
Program Repair: RepairBench (repairbench)
@@
Finally, we looked at the performance of LLM in automatic program repair with RepairBench (repairbench), which contained two benchmarks at the time: gitbugjava (90 tests, 23 variants) and defects4j (484 tests, 22 variants). A variant here is an LLM. Then, three versions of each benchmark were created based on the release dates of the LLMs. The first cut-off date is Oct. 2024 (9 variants left), and the second one is Jan. 2025 (17 variants left). Adding a final six additional benchmarks. The cost for these benchmarks is 1 for all tests, since the cost was global and not on a test basis.
@@ 6.3. RQ1: How many tests can be eliminated while maintaining the same ranking?
(a) All benchmarks
(b) Benchmarks with a cost reduction 39/50
Figure 3. Violin plot of the cost reduction (higher is better) of the different methods across different benchmarks with a target Kendall value of 1
In order to tackle this question, we will run our methods on all datasets targeting no ranking error. The aggregated results are plotted in a violin plot in Fig. 3, the full results are in LABEL:table:gain_full in the Appendix. The cost reduction is the fraction of cost saved, therefore higher is better.
For 11 of the benchmarks, no cost reduction was found, which means that for 39/50 we found a cost reduction. We observe that globally BISS outperforms the other methods. Interestingly the third quartiles of MILP and BISS are similar, meaning that they perform as well on the same fraction of benchmarks, however the worst case scenario for BISS is much better. MILP has a median at 0 cost reduction whereas BISS has a median nearly at 0.4. RANDOM offers better average and median case than MILP. However, it offers a worse third quartile; in other words, it often finds reductions, but they tend to not be as important as those found by MILP. RANDOM is on all aspects trailing behind BISS. GREEDY and PCA are clearly outperformed by the others.
When looking only at benchmarks where at least one method found a cost reduction, we clearly observe improved performances for all methods. However, these improvements do not change the relative performance of one method compared to another and BISS still outperforms the other methods with the same arguments.
We found that in practice not all solutions have perfect Kendall. This is due to our divide and conquer approach; without this component, the different methods have perfect ranking. So we introduce a metric the score combining both the cost reduction and the Kendall value. Intuitively, if the cost is reduced by 10%, for an optimal trade-off, the Kendall value should not be more than 10% lower. The score gives equal weight to a cost reduction and a Kendall value reduction: $score(c,k)=\frac{1}{2}(\frac{c}{2}+\frac{1+k}{2})$ where $c\in[0;1]$ is the fraction of cost reduced and $k\in[-1;1]$ is the Kendall coefficient. By default, if no reduction is made, the obtained score is 0.5; another way to think about this is that if the score is above 0.5, then the trade-off starts to become interesting. Therefore we plot in Fig. 4 the same data for the score and observe similar results as for the cost reduction.
(a) All benchmarks
(b) Benchmarks with a cost reduction 39/50
Figure 4. Violin plot of the score (higher is better) of the different methods across different benchmarks with a target Kendall value of 1
MILP by definition is the only method that guarantees a score of at least 0.5, which correspond to no cost reduction at all with the solution being the full test suite. BISS is still the method which offers the best results and rarely scores less than 0.5.
In more than 75% of the benchmarks, test instances were successfully removed with high Kendall. Reductions of costs range from a few percent up to almost 100% of the cost, with the best method averaging 44% cost reduction on all benchmarks and 55% on benchmarks where a reduction was found by any of the methods. BISS outperforms other methods. MILP and RANDOM then outperforms GREEDY and PCA by far.
(a) All benchmarks
(b) Benchmarks with a cost reduction 36/50
Figure 5. Violin plot of the cost reduction (higher is better) of the different methods across different benchmarks with a target Kendall of 0.99
For this RQ, we follow the same procedure as for RQ1, but with a Kendall target of 0.99, which allows for some leeway into the ranking while still being close to the original ranking. MILP is not present, because, as explained before (section 5), we could not adapt the MILP formulation to target a Kendall value different from 1 within the MILP framework.
The aggregated results are found in Fig. 5 and the per benchmark results are available in the Appendix in LABEL:table:gain_99_full. First, no cost reduction was found on 14 benchmarks, a worse result than for RQ1. This is explained by MILP being absent which was the only method to find a cost reduction on some benchmarks.
We observe the same trends as for Fig. 3. BISS outperforms the others, followed by RANDOM and GREEDY and PCA have similar performances. The mean cost reduction for BISS on all benchmarks is at 50% and nearly 70% on benchmarks where a cost reduction was found. This clearly highlights that reducing a bit the constraints by targeting a lower Kendall value can offer significant improvements in performances. As in RQ1, we als plot the same data for the score in Fig. 6.
(a) All benchmarks
(b) Benchmarks with a cost reduction 39/50
Figure 6. Violin plot of the score (higher is better) of the different methods across different benchmarks with a target Kendall value of 0.99
And we observe the same conclusions as in Fig. 4. BISS still outperforms all other methods. All other methods are hurt by the reduced target Kendall value and have more solutions with scores ¡ .5 than in RQ1 whereas it is not the case for BISS which has less of them.
The same relative performances trends are observed as in RQ1. BISS still outperform the others. BISS averages 50% cost reduction on all benchmarks and almost 70% on benchmarks where a reduction was found by any of the methods. Reducing a bit the target Kendall value enables leeway offering much larger reductions. Less benchmarks where found with a cost reduction despite the improvements because MILP was the only one to find improvements on these benchmarks.
(a) 25% of variants
(b) 50% of variants
(c) 75% of variants
(d) 100% of variants
Figure 7. Violin plot of the score (higher is better) of the different methods across all benchmarks with a target Kendall of 1 for different fraction of variants. The score is measured against the data with 100% of the variants. A score ¡ 0.5 is worse than not doing any minimization, this is indicated by the red dotted line which draws the limit.
For this RQ, we only consider MILP, BISS and RANDOM, the best performing methods for a target Kendall of 1. We sampled a fraction of all the variants for each benchmark, then ran the minimization procedure, and measured the cost reduction and Kendall on the full set of variants. Unlike RQ2, we only consider targeting a Kendall of 1, the same ranking, because we are already losing information by removing variants. This experiment reproduces the context of a competition using our approach. The benchmark is minimized with only a subset of all variants, then new variants are submitted by competitors, and the reduced benchmark is used. We chose to keep 75% 50%, and 25% of all variants. We sampled 10 random subsets for each percentage, with 10 seeds per subset and method, that is, 100 seeds in total per method.
However, there are two metrics to consider: the cost reduction and the Kendall. We measure the Kendall on the benchmark with 100% of the variants so the obtained Kendall may be much lower than the targeted one.
Since there are 50 different benchmarks, the supplementary material enables to plot individual data, here we will plot the aggregated data as violin plots of the score in Fig. 7. BISS generally outperforms MILP and RANDOM. Interestingly, when the number of variants is reduced the differences between BISS and RANDOM start to fade. That means, when the number of variants is reduced, it is more likely that less tests are necessary to fully rank them. In that case it is then easier to find subsets of tests that provide benefits; in other words, it becomes easier to sample them. When sampling good subsets of tests is easy, using BISS and RANDOM makes little difference, when it is harder the difference becomes important.
Overall, MILP rarely provides solutions with a score ¡ 0.5 compared to BISS and RANDOM; despite that its overall performance still trails behind BISS and RANDOM.
BISS still outperforms other methods despite the changes in variants. The less variants there are the more gains can easily be obtained. When few variants are used; it becomes easier to find solutions, differences between the performances of BISS and RANDOM fade.
6.6. RQ4: How many benchmark runs need to be made to make the minimization worth it?
(a) Target Kendall value : 1
(b) Target Kendallvalue 0.99
Figure 8. Violin plot of the break even points (lower is better) across the 12 benchmarks with costs for different target Kendall values with 100% of the variants. Break even points translate how many times the minimized benchmark needs to be run to offset the cost of the minimization. All values here are less than 1, indicating that after one run, cost savings have already been made.
To answer this RQ, we use the same data as in RQ1 and RQ2, we will compute the break even point. The break even point is the number of runs $n$ such that the cost of running the minimization is offset by the $n$ runs of the benchmark, the lower the better. More formally:
[…]
where $\mathcal{T}^{\prime}$ is the subset of tests returned by $M$ for the minimization of the test suite $\mathcal{T}$. This metric has no sense for benchmarks where no cost was defined therefore we exclude these, leaving us with 12 benchmarks. When no cost reduction was found, we replace the data point with the worst break even point among all methods among all 12 benchmarks, this avoids severely unbalanced data points and penalize methods for not finding any improvements. We plot the distributions on Fig. 8 for target Kendall of 1 and 0.99.. Technically, since we cannot run a benchmark partially, we can take the ceil of the break even point to get an integer number of minimized benchmark run; however that would leave no information on the Figure. Indeed, all our methods, even the baselines provide minimizations that are offset after less than 1 run of the minimized benchmark.
MILP is heavily penalized by the fact that most of the times it found no cost reduction and therefore is the worst performing method. BISS and RANDOM are the two best methods across the different target Kendall. RANDOM score points by being very fast compared to BISS despite finding worst cost reductions. Note that this does not take into account the obtained Kendall.
For all methods the cost of minimization is offset by less than one and at most two runs of the benchmark, highlighting the impact and potential savings of the RTSM problem. BISS and RANDOM outperform all other methods.
6.7. RQ5: How does our method BISS compare to the others?
  • BISS outperforms RANDOM with effect size of 0.52 and p-value ¡ $10^{-277}$
  • BISS outperforms MILP with effect size of 0.24 and p-value ¡ $10^{-308}$
  • RANDOM outperforms MILP with effect size of 0.22 and p-value ¡ $10^{-173}$
BISS outperforms all other methods; however, we note on a few benchmarks that using other approaches may lead to better solutions; more precisely, the MILP approach offers new perspectives for some benchmarks, whereas RANDOM fails on the same benchmarks as BISS.
6.8. RQ6: How does the different components affect performance?
Iterative solving
Timeouts with Divide and Conquer
In order to measure the impact of divide and conquer on the speed of our method, for all methods with divide and conquer on all different experiments we recorded the total number of timeouts. Then we looked for the number of timeouts of BISS and RANDOM without divide and conquer and iterative solving on the same scenario as RQ1, that is with all variants and with a Kendall of 1, this constitutes 1099 runs compared to the 58445 that we have with divide and conquer. As seen in Table 1, there is a large variation in the number of timeouts. Note that since MILP is not seed dependent, fewer runs are explaining the difference compared to the number from raw number to percentage.
Method BISS RANDOM PCA GREEDY MILP
Raw Number 144 40 5093 0 2650
Percentage 0.25% 0.07% 8.73% 0.00% 4.54%
Raw Number (No DC) 295 288 / / /
Percentage (No DC) 26.84% 26.21% / / /
7. Selected Case Studies
  1. (1)
    BNSL: Is the dataset highly redundant?
  2. (2)
  3. (3)
7.1. Case Study: BNSL: Is the dataset highly redundant?
Figure 10. Kendall (higher is better) with respect to cost ratio (left is better) for different methods for BNSL-2016 with a target Kendall value of 1
BNSL is a benchmark for score-based structure learning in Bayesian networks, the performance measured is the runtime. It comes from ASlib (bischl_aslib_2016) the algorithm selection collection of benchmarks. We chose to target BNSL since in (matricon21comparison), the authors find that it is easy to predict the best performing solver among a challenger and an incumbent using only a single test and emit the hypothesis that the benchmark is highly redundant. We want to confirm this hypothesis; therefore, we will look more precisely at the BNSL data extracted from our methods BISS and RANDOM the top performing ones on this scenario.
In Fig. 10 we plot the data points obtained for both methods and observe that all data points have very small cost ratios which basically amount to running a single test. BISS’ data points are agglomerated in the top left corner with a Kendall value fo 1 indicating that it is possible to have a perfect rankign with a single instance. Even RANDOM finds good reductions even though with more variance.
Overall, looking at the data, their benchmark is highly redundant as was hypothesized in (matricon21comparison).
BNSL, as hypothesized in (matricon21comparison), is a highly redundant dataset with a ranking of variants that can be done with only one test instance.
In this case study, we want to observe how the different distributional properties of the features of the tests change after minimization by BISS. We use the 10 different solutions with 100% of variants obtained with BISS from RQ1. Intuitively, the hope is that the minimization does not impact much the distributional properties of the features of the test suite such as min, max, etc. We use the features given in (bischl_aslib_2016) that were used for automated algorithm selection tasks. The protocol is the following: for each feature, we look at the relative change of distributional properties of the feature before and after minimization. The distributional properties selected are: min, 1st quartile, median, 3rd quartile, max, mean, and std. Since there are a lot of different features, we compute the median relative change across all features across all seeds. Therefore for each distributional property we have a list of median relative change, one per seed, which we again aggregate by min, median, mean and max on the rows. Then in Table 2 and Table 3, we show this evolution for SAT18 and SAT20.
For SAT18, a lot of test are easily removed, at least 38%, but usually more than 75%. Despite that, the relative change of the other distributional properties are all less than 11%, except for Q3 and max. It is to be expected, since we want to minimize the cost, and instances with large features are often the instances, which take the most time therefore are more likely to be removed, therefore affecting max and Q3. For SAT20, some tests are easily removed, at least 24% and usually more than 30%, despite that, the relative change of the other distributional properties are all less than 9%. On both benchmarks, the relative change in features is minor compared to the changes made in the test suite.
Table 2. Min, median, mean and max of relative median change across all features for solutions found by BISS on SAT 18
% of tests kept mean std min Q1 median Q3 max
min 38.81% 4.38% 4.26% 0.00% 3.25% 5.77% 7.03% 0.00%
median 75.78% 8.71% 7.62% 0.00% 4.72% 7.60% 16.16% 6.95%
mean 68.22% 8.39% 7.34% 0.82% 4.94% 7.71% 17.88% 9.01%
max 84.99% 12.98% 10.59% 8.18% 8.50% 10.33% 36.87% 26.43%
Table 3. Min, median, mean and max of relative median change across all features for solutions found by BISS on SAT 20
% of tests kept mean std min Q1 median Q3 max
min 24.41% 3.62% 3.13% 0.00% 0.67% 1.17% 1.60% 0.00%
median 30.74% 4.12% 4.40% 0.00% 1.51% 1.93% 3.04% 0.00%
mean 37.51% 4.85% 5.05% 0.00% 1.87% 2.70% 3.91% 1.11%
max 71.38% 8.60% 8.92% 0.00% 3.58% 7.98% 8.81% 6.97%
(a) cutoff on 2024/10/01 (9 variants/LLMs)
(b) cutoff on 2025/01/01 (17 variants/LLMs)
(c) no cutoff (22 variants/LLMs)
RepairBench from Silva and Montperrus (repairbench) is based on Defects4J (defects4j) to measure the program repair capabilities of LLMs. To evaluate the evolution of the skills of LLMs, here, the variants, we plot on Fig. 11 the different data points of Kendall values with respect to the cost ratio for our methods. The Kendall value is measured on the full dataset, that is, with all LLMs. Here, the cost ratio is the ratio of tests kept in the minimized benchmark.
Interestingly, for the cutoff date in 2024, we observe a very low cost ratio, indicating that most LLMs before this date can be easily ranked. The Kendall value, however, is quite low, indicating that the behavior of the more recent LLMs cannot be correctly captured by the present minimized test set.
For the cutoff date in 2025, we observe that the Kendall value jumps significantly, translating the fact there has been a lot of progress of LLMs on the task. Note that at that date, there were no ‘reasoning’ LLMs; this is why we also observe that the Kendall value is not yet close to 1, translating that the reasoning behavior is not captured by the previous models. In terms of cost ratio, we observe heterogeneous results that are actually quite similar to the cost ratio distribution of the case with no cutoffs.
Finally, with no cutoffs, we observe that we can have cost reductions of more than 90% of the tests set with no loss of accuracy in the best case, indicating that the benchmark is quite redundant for measuring the performances difference among the LLMs considered. With MILP, we obtain a cost ratio of 5%, meaning that all the LLMs’ performances can be correctly ranked with only 5% of the tests, which can save hundreds of dollars per attempt according to the costs reported by (repairbench).
Defects4J contains a lot of redundant instances that bring little to no discriminative power to rank the different LLMs. Indeed, 5% of the 484 bugs are enough to perfectly rank all the LLMs, saving hundreds of dollars per attempt. Interestingly, RepairBench has also been discontinued accord to their website [^1] https://repairbench.github.io/ last consulted Dec. 2 2025. The evolution of the cutoff dates and the results show that LLMs before Oct. 2024 behaved quite differently from the ones after, the LLMs released between Oct. 2024 and Jan. 2025 are more similar in terms of performance to the ones released afterwards.
8. Threats to Validity
Internal threats.
Timeouts
Stability
Generalizability
9. Discussion and Conclusion
Recommendations for software developers. Developers consume benchmarks for performance comparison, nonregression, or tuning. Smaller benchmarks with the same discriminative power mean that consumers will maintain accuracy while spending less time executing their software on variants. For scenarios like tuning or nonregression, where benchmarks are run repeatedly, the gains are amplified with each run, resulting in substantial overall benefits. We thus offer the same recommendations as for benchmark design, and empirical results show that important gains can be achieved for reducing the computational cost.
Cost saving potentials. A highlight of our results is, for example, defects4j, which can be reduced to only 5% of the tests, saving up to hundreds of dollars per new LLM test. These reductions can be observed on many different datasets at different scales. Our results are available in the supplementary material, and since our tool is available, this enables benchmark creators to easily check their benchmark.
Conclusion. We have introduced the ranked test suite minimization (RTSM) and its solution. It is a new problem that aims to minimize a benchmark while keeping the performance ranking of the variants over the test suites intact. This problem is highly relevant for practitioners, as it enables cost-effective benchmarking for variant comparison, rapid regression detection, performance modeling, benchmark design, etc.
In general, with BISS, more than half of the benchmarks could be reduced with a mean reduction over these reducible benchmarks of 67%. This implies that quite extensive reductions are available for many benchmarks, including those used today.
Future Work. Next steps in the direction of this work include studying the applications and implications of our methods in different domains. For example, the field of AI is driven by benchmark, standardizing such minimization approaches could drastically reduce the cost of research comparison, thus increasing research velocity. An important direction is studying other formulations of the linear regression problem in order to make the divide and conquer algorithm always return solutions with the guaranteed target Kendall value. Another direction is to work with partial data, that is, with some variants running only on a subset of the tests.
Appendix A Full Data